# What Lurks Within? Concept Auditing for Shared Diffusion Models at Scale

### v1

@@ 6.1. Statistic Analysis on Civitai
We first collect and analyze meta from $100,200$ LoRA DMs available on Civitai at the time of collection.
Our analysis reveals that most DMs do not specify trigger words in their metadata (Figure 10). In some cases, “unclaimed” prompts like “ohwx,” “sks” are used. These prompts, which are not part of the model vocabulary, are often introduced during fine-tuning to trigger unique concepts. However, the absence of trigger words or the use of uncommon words creates significant challenges for concept auditing, as the intended concepts cannot be easily inferred from the metadata.
To perform a comprehensive evaluation, we conduct an analysis on the models hosted by Civitai. Based on the analysis result (see Section D in Appendix for details), we identify four major topics of concepts: “celebrity,” “game,” “cartoon,” and “movie,” and a dominant DM: Stable Diffusion 1.5 (SD1.5). Hence, we download and analyze the LoRA models from these four categories and with Stable Diffusion 1.5 as the base model.
Figure 10. Common trigger words on Civitai. N/A indicates that the user does not provide trigger words.
@@ 6.2. Model Collection
The model collection process is as follows. First, we search for the LoRA models using the tags provided by each model in Civitai. For the celebrity topic, we use the tag “celebrity,” “actress,” “real_person,” and “actor.” For the cartoon topic, we use the tag “cartoon,” “manga,” and “anime.” For the videogame topic, we use the tag of “videogame,” “videogames,” and “video_game.” For the movie topic, we use the tag of “movie.” Second, we download the most frequently downloaded LoRA models in the search results. Third, since the model files may stored in different formats (e.g., different parameter names), we standardize all DMs with consistent formats. Lastly, we manually remove DMs with duplicate concepts (e.g., multiple DMs fine-tuned on the same celebrity or character) and exclude style-related LoRA models that are not concept-specific.
After filtering, we have collected $174$ celebrity models, $145$ cartoon models, $192$ videogame models, and $179$ movie models.
@@ 8.1. Scope of Concepts
Defining what constitutes a “concept” in DMs remains an open and underexplored problem. Depending on the context, a concept may refer to a broad category, such as “Disney characters,” or a highly specific instance, like “Mickey Mouse.” This ambiguity exists across related areas such as concept customization (kumari2023multi, ; zeng2024improved, ; smith2023continual, ) and concept erasure research (Gandikota2023Erasing, ; lu2024mace, ; liu2024implicit, ), where there is no consistent standard for what counts as a distinct or meaningful concept. This ambiguity presents a challenge for concept auditing, as it is difficult to define universal criteria.
In this paper, rather than relying on a rigid definition, we adopt an example-based approach: a concept is considered present if the model can generate recognizable outputs that are aligned with a small set of reference examples. This flexible design enables our framework to handle a broad range of concept granularities and types. In our evaluation, we include concepts such as individual celebrities, full character sets from specific cartoons (e.g., all characters from Pokémon), and object classes from video games. We leave further exploration of formal concept taxonomies and boundaries as future work, particularly toward improving auditing resolution across different concept types.
@@ 8.3. Toward Adversarial Auditing
This work focuses on non-adversarial auditing, where fine-tuned DMs are not deliberately optimized to evade auditing. This setting reflects the majority of real-world cases on public platforms, where users unknowingly or negligently publish potentially problematic DMs, often without consistent metadata or clear prompt disclosure. To assess robustness under adversarial conditions, we include a controlled experiment with a simple backdoor-style trigger (Tables 3 and 4), where a concept is activated by an unrelated and abnormal prompt. PAIA remains effective in this setting, highlighting the robustness of its model-centric design.
@@ 9. Conclusion
We introduce Prompt-Agnostic Image-Free Detection (PAIA), a model-centric framework for auditing fine-tuned diffusion models without relying on prompt optimization or image-based detection. PAIA combines two key innovations: a prompt-agnostic design that analyzes internal model behavior during prompt-insensitive stages of generation, and an image-free mechanism based on conditional calibrated error, which compares denoising behavior against the base model to reveal concept learning.
Extensive experiments demonstrate that PAIA consistently outperforms state-of-the-art baselines in both accuracy and efficiency. On $320$ benchmark models, PAIA achieves accurate and efficient concept auditing with significantly reduced computation. On $690$ real-world models collected from Civitai, it reaches an average accuracy of over $92$% across categories including celebrities, cartoons, video games, and movies, highlighting its robustness and real-world applicability.
@@ Appendix A Proof of Lemma 1
where $\delta_{jk}$ is the Kronecker delta, $\delta_{jk}=1$ if $j=k$ and 00 otherwise. For simplicity, we denote the Jacobian of $S_{i}$ with respect to $A$ as:
@@
Appendix B Terms used in random prompt generation.
@@
Appendix C Benchmarking Dataset Samples: Celebrities and Cartoons
We provide example images from the curated datasets used in our benchmark evaluation. Each concept is represented by a small set of reference images used during fine-tuning and for auditing evaluation.
@@ Appendix C Benchmarking Dataset Samples: Celebrities and Cartoons
Appendix D Civitai model analysis.
@@
Appendix E Trigger Quality Analysis on Civitai Models

### v2

@@ 5.1. Statistic Analysis on Civitai
We first collect and analyze metadata from $100,200$ LoRA DMs available on Civitai at the time of collection. Our analysis reveals that most DMs do not specify trigger words in their metadata (Figure 12). In some cases, “unclaimed” prompts like “ohwx,” “sks” are used. These prompts, which are not part of the model vocabulary, are often introduced during fine-tuning to trigger unique concepts. However, the absence of trigger words or the use of uncommon words creates significant challenges for concept auditing, as the intended concepts cannot be easily inferred from the metadata.
To perform a comprehensive evaluation, we conduct an analysis on the models hosted by Civitai. Based on the analysis result (Appendix Section E), we identify four major topics of concepts: “celebrity,” “game,” “cartoon,” and “movie,” and a dominant DM: Stable Diffusion 1.5 (SD1.5). Hence, we download and analyze the LoRA models from these four categories and with Stable Diffusion 1.5 as the base model.
Figure 12. Frequency of common trigger words on Civitai. N/A indicates that the user does not provide trigger words.
@@ 5.2. Model Collection
The model collection process is as follows. First, we search for the LoRA models using the tags provided by each model in Civitai. For the celebrity topic, we use the tag “celebrity,” “actress,” “real_person,” and “actor.” For the cartoon topic, we use the tag “cartoon,” “manga,” and “anime.” For the videogame topic, we use the tag of “videogame,” “videogames,” and “video_game.” For the movie topic, we use the tag of “movie.” Second, we download the most frequently downloaded LoRA models in the search results. Third, since the model files may stored in different formats (e.g., different parameter names), we standardize all DMs to a consistent format before performing analysis. Lastly, we manually remove DMs with duplicate concepts (e.g., multiple DMs fine-tuned on the same celebrity or character) and exclude style-related LoRA models that are not concept-specific. After filtering, we have collected $174$ celebrity models, $145$ cartoon models, $192$ videogame models, and $179$ movie models.
@@ 5.3. Auditing Performance of PAIA
5.4. Auditing Performance on Rare Concepts
Table 9. Auditing performance of PAIA on Civitai DMs with rare concepts.
Category Common concepts Rare concepts
Celebrity 93.97% 96.43%
Cartoon 96.90% 93.75%
Videogame 95.83% 95.00%
Movie 92.74% 96.15%
Previous analysis explores PAIA’s performance on the common concepts with most downloads. For rigorous evaluation, we further assess PAIA’s performance on rare concepts (e.g., less popular identities, new characters). We collected Civitai LoRA models with fewer than $200$ downloads and manually filter out models with common concepts. In total, $81$ Civitai LoRA models have been collected, about $20$ per category. The outlier detector was trained on common concepts and tested on these rare ones. As shown in Table 9, PAIA achieves comparable or even better performance. This results show that PAIA can be well generalized to rare concepts.
@@ 7.1. Scope of Concepts
Defining what constitutes a “concept” in diffusion models remains an open and underexplored problem. Depending on context, a concept may refer to a broad category, such as “Disney characters,” or a highly specific instance, like “Mickey Mouse.” This ambiguity is well-documented in related areas such as concept customization (kumari2023multi, ; zeng2024improved, ; smith2023continual, ) and concept erasure (Gandikota2023Erasing, ; lu2024mace, ; liu2024implicit, ), where no consistent standard exists for what qualifies as a distinct or meaningful concept. Such ambiguity poses a challenge for concept auditing, as it complicates the definition of universal auditing criteria.
Rather than imposing a rigid definition, we adopt an example-based approach: a concept is considered present if the model can generate recognizable outputs aligned with a small set of reference examples. This flexible design enables our framework to support a wide range of concept granularities and types. In our evaluation, we audit models on concepts such as individual celebrities, complete character sets from specific cartoons (e.g., all Pokémon characters), and object categories from video games. We leave the development of formal taxonomies and clearer conceptual boundaries as an important direction for future work, particularly for improving auditing resolution across different semantic scopes.
While our focus is on semantically meaningful and legally actionable concepts, we note that concept-level auditing for unsafe content (e.g., NSFW or violent imagery) introduces additional ambiguity. Safety-related analysis of diffusion models has gained increasing attention (schramowski2023safe, ; zhang2024generate, ; yang2024sneakyprompt, ; Li2024SafeGenMU, ; han2024shielddiff, ; chen2025safe, ; li2025detect, ; chen2025comprehensive, ), yet defining and auditing “unsafe content” remains difficult due to its context-dependence, vague definitions, and heterogeneous semantics. These factors make unsafe content less applicable to our example-based formulation. While PAIA is not designed to audit unsafe content directly, its methodology could potentially be extended with domain-specific knowledge or safety taxonomies in future work.
@@ 7.3. Toward Adversarial Auditing
This work focuses on non-adversarial auditing, where fine-tuned DMs are not deliberately optimized to evade auditing. This setting reflects the majority of real-world cases on public platforms, where users unknowingly or negligently publish potentially problematic DMs, often without consistent metadata or clear prompt disclosure. To assess robustness under adversarial conditions, we include five adaptive attacks where the attacker has full knowledge of the auditing framework. PAIA remains effective in both settings, highlighting the robustness of its model-centric design.
@@ 8. Conclusion
We introduce Prompt-Agnostic Image-Free Auditing (PAIA), a model-centric framework for auditing fine-tuned diffusion models without relying on prompt optimization or image-based detection. PAIA combines two key innovations: a prompt-agnostic design that analyzes internal model behavior during prompt-insensitive stages of generation, and an image-free mechanism based on conditional calibrated error, which compares denoising behavior against the base model to reveal concept learning.
Extensive experiments demonstrate that PAIA consistently outperforms state-of-the-art baselines in both accuracy and efficiency. Our auditing setting assumes internal access to DMs, but does not require access to proprietary fine-tuning data or user prompts, an assumption aligned with how hosted platforms audit uploaded models. On benchmark models, PAIA achieves accurate and efficient concept auditing with significantly reduced computation. On real-world models collected from Civitai, it reaches an average accuracy of over 9292% across categories including celebrities, cartoons, video games, and movies. In addition, PAIA is robust against three adaptive attacks where the attackers know the auditing design, highlighting its robustness and real-world applicability.
@@
9. Acknowledgments
The work of X. Yuan and L. Zhang was supported in part by the National Science Foundation under Grant 2427316, 2426318. The work of L. Guo was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-24-1-0044. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Proof of Lemma 1
where $\delta_{jk}$ is the Kronecker delta, $\delta_{jk}=1$ if $j=k$ and 0 otherwise. For simplicity, we denote the Jacobian of $S_{i}$ with respect to $A$ as:
@@
Appendix B PAIA Pseudocode
We present a complete pseudocode of PAIA in Algorithm 1.
Input: Target model $f_{W^{\prime}}$, Base model $f_{W}$, Image set with the target concept $\mathcal{D}_{\text{target}}$, Image set with irrelevant concepts $\mathcal{D}_{\text{irr}}$, Cutoff time step $\tau$, Time steps $\mathcal{T}$
Output: Prediction $y\in\{0,1\}$ indicating whether $f_{W^{\prime}}$ can generate the target concept
1
2 Freeze the cross-attention in $f_{W^{\prime}}$ to obtain parameters $W^{\prime\prime}$
3 Function ComputeFeatures($f_{W^{\prime}}$, $\mathcal{D}$):
4   Initialize feature set $\mathcal{F}\leftarrow\emptyset$
5 foreach $\bm{x}\in\mathcal{D}$ do
6    foreach $t\in\mathcal{T}$ do
7         Generate noised latent image representation $\mathbf{z}_{t}$ and pseudo-prompt $\mathbf{p}$ from $\bm{x}$
8       if $t<\tau$ then
9          $\mathcal{L}_{cce}^{t}\leftarrow\mathbb{E}\left[\|{\bm{\epsilon}}_{W^{\prime}}(\mathbf{z}_{t},\mathbf{p})-\bm{\epsilon}_{0}\|^{2}\right]-\|{\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})-\bm{\epsilon}_{0}\|^{2}$
10       else
11          $\mathcal{L}_{cce}^{t}\leftarrow\mathbb{E}\left[\|{\bm{\epsilon}}_{W^{\prime\prime}}(\mathbf{z}_{t},\mathbf{p})-\bm{\epsilon}_{0}\|^{2}\right]-\|{\bm{\epsilon}}_{W}(\mathbf{z}_{t},\mathbf{p})-\bm{\epsilon}_{0}\|^{2}$
12
13    $\phi(\bm{x})\leftarrow\{\mathcal{L}_{cce}^{t}\}_{t\in\mathcal{T}}$
14    $\mathcal{F}\leftarrow\mathcal{F}\cup\{\phi(\bm{x})\}$
15 return $\mathcal{F}$
16
17 Training Phase:
18 $\mathcal{F}_{\text{irr}}\leftarrow$ ComputeFeatures($f_{W^{\prime}}$, $\mathcal{D}_{\text{irr}}$)
19 Train an anomaly detector $\mathcal{A}$ on $\mathcal{F}_{\text{irr}}$
20 Inference Phase:
21 $\mathcal{F}_{\text{target}}\leftarrow$ ComputeFeatures($f_{W^{\prime}}$, $\mathcal{D}_{\text{target}}$)
22 Initialize anomaly indicator list $Y\leftarrow\emptyset$
23 foreach $\phi(\bm{x}_{i})\in\mathcal{F}_{\text{target}}$ do
24   Compute anomaly score $s_{i}\leftarrow\mathcal{A}(\phi(\bm{x}_{i}))$
25 $y_{i}\leftarrow\mathbb{I}[s_{i}>\delta]$ // 1 if the anomaly score is great than threshold $\delta$
26 $Y\leftarrow Y\cup\{y_{i}\}$
return $y=\text{MajorityVote}(Y)$
Algorithm 1 PAIA: Prompt-Agnostic Image-Free Auditing (Training and Inference)
Appendix C Terms used in random prompt generation.
@@ 26 $Y\leftarrow Y\cup\{y_{i}\}$
Appendix D Benchmarking Dataset Samples: Celebrities and Cartoons
We provide example images from the curated datasets used in our benchmark evaluation in Figure 13 and 14. Each concept is represented by a small set of reference images used during fine-tuning and for auditing evaluation.
@@ Appendix D Benchmarking Dataset Samples: Celebrities and Cartoons
Appendix E Civitai model analysis.
@@
Appendix F Additional Evaluation Results
F.1. Impact of Cutoff Time Step.
Figure 17. Impact of cutoff time step.
We evaluate the impact of varying the cutoff timestep $\gamma$ in CCE (Eq. 11) using Civitai models, with $\gamma\in\{\frac{T}{8},\frac{T}{4},\frac{3T}{8},\frac{T}{2},\frac{5T}{8},\frac{3T}{4},\frac{7T}{8},T\}$. As shown in Figure 17, the best performance is achieved around $\gamma=\frac{T}{2}$, balancing prompt suppression and behavioral signal strength. Early cutoffs underutilize prompt information, while late cutoffs weaken calibration. Performance across mid-range values (e.g., between $\frac{3T}{8}$ and $\frac{5T}{8}$) is relatively stable , indicating that PAIA is not sensitive to the exact choice of $\gamma$.
F.2. Effectiveness of Conditional Calibration on Civitai Models
Table 11. Impact of Conditional Calibration on Civitai Models.
Category With Conditional Calibration Without Conditional Calibration
Celebrity 93.97% 91.38%
Cartoon 96.90% 95.86%
Videogame 95.83% 95.05%
Movie 92.74% 91.06%
We investigate the effectiveness of conditional calibration on Civitai Models. As shown in Table 11, conditional calibration improves auditing accuracy across all categories. Although the gains are relatively small (1–2%), they are consistent and meaningful, especially given the strong baseline performance without calibration.
Appendix G Trigger Quality Analysis on Civitai Models